OpenStudy (anonymous):

MEDALS WILL BE GIVEN The ordered pairs (1, 4) (2, 16) (3, 64) and (4, 256) represent a function. Write an equation that represents this func…

OpenStudy (anonymous):

the answers

OpenStudy (anonymous):

there is no answers you have to figure out the questions

OpenStudy (anonymous):

functions

OpenStudy (anonymous):

4=4/1 16=4/2 64=4/3

OpenStudy (anonymous):

he's a magical wizard *-*

OpenStudy (anonymous):

EAT CACA

OpenStudy (anonymous):

rofl

OpenStudy (anonymous):

the answer is b)y = 4^x

OpenStudy (anonymous):

your suppose to write an equation there is no answer. smart.